# Meta4XNLI: A Crosslingual Parallel Corpus for Metaphor Detection and Interpretation

### v1

@@ 5 Results
Multilingual: In this set of experiments we assembled Meta4XNLIEN and Meta4XNLIES to train the MLMs. The evaluation is conducted for each language separately and results detailed in Table 9. Best results in ES are obtained by mDeBERTa, which are higher than the top result from monolingual experiments. In EN, mDeBERTa is the model that achieves better performance but very close to that of XLM-RoBERTa. The highest F1 score is 8 points lower than that of ES but outweighs EN monolingual results. This suggests that the combination of parallel multilingual data for training is beneficial for the performance of the models.
| mDeBERTa | XLM-RoBERTa | ||||||
| Train | Test | F1 | Inv. | Oov | F1 | Inv | Oov |
| ES + EN | ES | 68.70 1.27 | 72.63 1.23 | 63.03 1.95 | 66.01 0.55 | 70.29 1.18 | 59.84 1.43 |
| EN | 60.80 1.05 | 63.16 1.38 | 55.12 1.08 | 60.14 1.32 | 61.57 1.14 | 56.79 1.88 | |
@@
Zero-shot cross-lingual: In these experiments we perform evaluation of Meta4XNLI in the opposite language to that utilised for training. Results are reported in Table 10. XLM-RoBERTa performance exceeds that of mDeBERTa in both languages. Nonetheless, F1 score for EN is almost 20 points lower than ES evaluation results. In addition to the differences in annotation criteria between languages and datasets, another aspect to bear in mind in this scenario could be the number of positive examples present in training sets. As we explained in Section 3.3, Meta4XNLIEN contains a higher number of metaphorical instances, thus models are exposed to a greater variety of examples that can be transferred to ES. While a more reduced amount of instances in Meta4XNLIES seen during training might hinder model’s generalization ability, as the low recall and high precision scores show in Table 19 in Appendix B.
@@
In the first setup, we fine-tuned the models for the NLI task with MultiNLI dataset Williams, Nangia, and Bowman (2018). Then we conducted the evaluation with two different splits for each source dataset conforming Meta4XNLI: pairs with at least one metaphorical expression and pairs lacking metaphors. From accuracy scores reported in Table 11 we can observe certain variability in the results. In the majority of cases, XLM-RoBERTa achieves better performance than mDeBERTa. We do observe a tendency of higher results on the sets of pairs without metaphors than on the set with metaphors. The exceptions to this current are the results from XLM-RoBERTa for XNLIdev and XNLItest in ES. In these partitions, the subset of pairs with metaphorical expressions obtained better results, although the difference does not even reach one point. We hypothesize the original language of the dataset might be involved in this performance, since XNLI includes natural utterances of EN that were afterwards manually translated to ES, thus some artifacts might have been introduced during this process Artetxe, Labaka, and Agirre (2020).
The second scenario consisted in fine-tuning the models on two setups: a) with pairs without metaphors and b) pairs with and without metaphors. We performed the evaluation on subsets split by the criterion of metaphor occurrence as well. Results reported in Table 12 show XLM-RoBERTa outperforming mDeBERTa in all contexts. Both models show best results for the NLI task on the examples without metaphors and the lowest performance with pairs that contain metaphorical expressions. This outcome replicates in both languages and experimental setups a) and b).
@@ 6 Error Analysis
In this section, we manually inspected a subset of erroneous cases in order to provide a qualitative insight of results and with the intention of finding potential explanations of errors and models performance for both detection and interpretation tasks.
@@
We selected the predictions from the model that obtained highest F1 score from the monolingual experiments in ES for both Meta4XNLI and CoMeta evaluations. We extracted false negatives and false positives and grouped the tokens by their number of occurrences. Within the false positives of CoMeta test set, we find tokens like apoyo (lit. “support”) or tensión (lit. “tension”) that appear more frequently used figuratively than with their literal sense or even only appear in the training set labeled as metaphor. Other wrong predictions are words that appeared in specific domains, such as texts that allude to the pandemic, e.g. ola (lit. “wave”), not detected as metaphorical due to its absence with metaphorical meaning in Meta4XNLI sentences. Something similar occurs with the misclassified tokens from Meta4XNLI test set. Most errors stem from conventional metaphors, namely gran, abrir, paso, claro (lit. “great”, “to open”, “step”, “clear”) that occur regularly used with their metaphorical meaning. The lack of balanced examples might contribute to these predictions. However, we maintained the distribution as is, since our aim is to study the presence and prevalence of metaphor in natural language utterances.
@@
We analysed a subset of 30 errors from each experimental setup, both EN and ES, and evaluation sets. We chose the predictions from XLM-RoBERTa since it is the model that performs better in this task in most scenarios. Although results show that models struggle more to identify the inference relation if there is a metaphorical expression involved in the pair sentences, we do not observe any particular feature within the errors. A remarkable aspect to highlight is the lower results of esXNLI with respect to XNLIdev and XNLItest in Table 11. It could be motivated by the difference of the text domains, since XNLI is an extension of MultiNLI, which maintains the same set of textual domains. While esXNLI is a collection of texts from another set of genres and sources. Regarding EN, some of the errors might derive from the misclassification of some pairs, since annotations were developed on ES text. A metaphorical expression involved in the inference relationship in ES might not be present in its EN version and vice versa. Thus, samples from these two classes, pairs with and without metaphors, should be reexamined in EN and correctly classified for future experimentation.
@@ 7 Conclusions and Future Work
In this work, we provided Meta4XNLI, a cross-lingual parallel dataset in ES and EN labeled for metaphor detection and interpretation framed within the task of NLI. With this new resource that contains parallel text and annotations, we developed a series of experiments to assess the capabilities of MLMs when dealing with this kind of figurative expressions present in natural language utterances.
@@
Regarding the task of metaphor detection, after the annotation process and experiments results, we can conclude the importance of establishing a unified criterion for annotation that is valid for different languages if the aim is to continue researching cross-lingual approaches. In addition, the semi-automatic process of annotation followed for EN shows that automatic labeling of cross-lingual metaphor is far from trivial. Metaphorical expressions are language and culture dependant. Moreover, the translation of the data introduces a new layer in which metaphorical expressions can either be lost from the translation of the source to the target language, or can be introduced in the target language by means of the translator, either human or automatic. Further work to explore automatic annotation methodologies would be of considerable value in order to reduce the demanding workload and effort of manual labelling in more than one language.
@@
With respect to results, the purpose of our experiments is not to overcome the state-of-the-art results, but to evaluate models performance from a cross-lingual approach. Best results are obtained when Meta4XNLI in both languages is used for training. The augmentation of the training set size and the parallel annotations might boost this performance. Cross-domain and monolingual experiments show how the lack of consistency in the annotation criteria affects the performance of models. This can be observed as well in the zero-shot cross-lingual setup, although the scenario of training in EN and evaluating in ES shows competitive performance. It should be noted that the EN set contains a larger number of instances annotated as metaphorical in the training set. In addition, the in-vocabulary and out-of-vocabulary evaluation points to some kind of bias in the learning process. This could stem from the fact that the majority of the metaphor instances are conventional or due to lexical memorization Levy et al. (2015); Boisson, Espinosa-Anke, and Camacho-Collados (2023). Future research on this line of work should be carried out to clarify this issue.
@@
For metaphor interpretation, we evaluated the ability of MLMs to understand metaphorical expressions framed within the task of NLI. We provide parallel annotations at premise-hypothesis pair level that mark if the presence of metaphorical expressions is relevant for the inference relationship. We exploited this information to conduct our experiments. From reported results, we can recognize a tendency of the models to show a lower performance with pairs that contain at least one metaphorical expression. However, this trend breaks with XLM-RoBERTa results when fine-tuned in ES and evaluated on the XNLIdev and XNLItest in the same language. Since these sets are sentences translated from EN, we presume the translation process might induce biases in metaphor occurrence and the “naturalness” of the sentences. Similar to metaphor detection, future work to analyse the impact of translation in the development of metaphor parallel resources should be explored for the task of metaphor interpretation, as well as additional experimentation from a multilingual perspective.
@@ 8 Limitations
Metaphor annotation is an inherently subjective task. This variance in annotations is reflected on Meta4XNLIEN, due to the different criterion employed through the annotation process. Labels in this language should be updated and further revised to improve their quality. Disagreement and subjectivity could be counterbalanced by a larger number of annotators, in order to develop more consistent and reliable labeled data, however it requires a huge amount of time and workload. Data augmentation and semi-automatic methods could be exploited to create larger datasets with similar characteristics to the one we present and extend it to more languages, since most corpora available for metaphor processing is of reduced size and limited to a narrow set of languages. The existence of parallel resources in multiple languages others than EN that reflect cultural and real world knowledge nuances is of great importance to continue researching such a complex phenomenon as figurative language, specifically metaphors.
Acknowledgements.
This work has been supported by the HiTZ center and the Basque Government (Research group funding IT-1805-22). Elisa Sanchez-Bayona is funded by a UPV/EHU grant “Formación de Personal Investigador”. We also thank the funding from the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and ERDF A way of making Europe; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR. Rodrigo Agerri currently holds the RYC-2017-23647 (MCIN/AEI/10.13039/501100011033 and ESF Investing in your future) fellowship.References
[…]
- Leong, Beigman Klebanov, and Shutova (2018) Leong, Chee Wee (Ben), Beata Beigman Klebanov, and Ekaterina Shutova. 2018. A Report on the 2018 VUA Metaphor Detection Shared Task. In Proceedings of the Workshop on Figurative Language Processing, pages 56–66, Association for Computational Linguistics, New Orleans, Louisiana.
- Raffel et al. (2023) Raffel, Colin, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. 2023. Exploring the limits of transfer learning with a unified text-to-text transformer.
Appendix A Data Splits Statistics
@@
Appendix B Detection Experiments Results
@@
| mDeBERTa | XLM-RoBERTa | ||||||
| Train | Test | F1 | Prec. | Rec. | F1 | Prec. | Rec. |
| ES + EN | ES | 68.70 1.27 | 72.78 3.45 | 65.29 3.77 | 66.01 0.55 | 70.35 4.42 | 62.53 3.93 |
| EN | 60.80 1.05 | 65.03 2.08 | 57.23 3.10 | 60.14 1.32 | 64.97 2.81 | 56.15 3.04 | |

### v2

@@ 5 Results
| Llama-3.1-8B-Instruct | Qwen2.5-7B-Instruct | gemma-7b-it | ||||||||
| Train | Test | F1 | Inv | Oov | F1 | Inv | Oov | F1 | Inv | Oov |
| M4XES | M4XES | 59.15 | 63.77 | 53.03 | 56.76 | 62.86 | 48.61 | 58.06 | 63.33 | 51.17 |
| CoMeta | 50.06 | 56.55 | 42.74 | 44.26 | 53.79 | 33.23 | 45.32 | 57.78 | 36.42 | |
| M4XEN | M4XEN | 50.41 | 53.58 | 44.57 | 47.66 | 51.37 | 41.63 | 52.61 | 56.37 | 45.97 |
| VUAM | 26.78 | 22.31 | 38.47 | 26.44 | 21.37 | 40.39 | 26.72 | 22.13 | 38.94 | |
In both languages, in-vocabulary evaluation results are higher than the overall ones and out-of-vocabulary results, lower. This is not the case when we use VUAM corpus for testing the model fine-tuned with Meta4XNLIEN. In this setup, similarly to results from cross-domain experiments, performance drastically falls around 30 points of F1 score and out-of-vocabulary results are the highest. The reason might be the mismatches in the annotation process. In ES, when evaluating CoMeta, we also encounter a decrease in the performance. However, it is a smaller difference of 10 points. This might be caused by the variety of text genres and dissimilarities between sentences from each dataset. Overall, encoder-only models outperform decoder-only models in this task. Despite the decrease of F1 scores, we can observe the same tendencies as with MLM in the results of LLMs in Table 12.
Multilingual: In this set of experiments, we assembled Meta4XNLIEN and Meta4XNLIES to train MLMs and LLMs. The evaluation is conducted for each language separately, and the results are detailed in Table 13. Encoder-only models outperform decoder-only models in all cases, as in the monolingual experimental setup. Best results in ES are obtained by mDeBERTa and Llama-3.1-8B-Instruct, which are higher than the top result from monolingual experiments. In EN, mDeBERTa is the model that achieves better performance but is very close to that of XLM-RoBERTa. Llama-3.1-8B-Instruct outperform Qwen2.5-7B-Instruct and gemma-7b-it by a larger margin. The highest F1 score is 8 points lower than that of ES but outweighs EN monolingual results. This suggests that the combination of parallel multilingual data for training is beneficial for the performance of the models.
| Meta4XNLIES | Meta4XNLIEN | |||||
| F1 | Inv | Oov | F1 | Inv | Oov | |
| mDeBERTa | 68.70 1.27 | 72.63 1.23 | 63.03 1.95 | 60.80 1.05 | 63.16 1.38 | 55.12 1.08 |
| XLM-RoBERTa | 66.01 0.55 | 70.29 1.18 | 59.84 1.43 | 60.14 1.32 | 61.57 1.14 | 56.79 1.88 |
| Llama-3.1-8B-Instruct | 61.46 | 64.52 | 57.40 | 55.26 | 57.91 | 50.00 |
| Qwen2.5-7B-Instruct | 58.72 | 63.35 | 52.60 | 52.62 | 56.56 | 46.25 |
| gemma-7b-it | 57.02 | 64.38 | 47.08 | 54.60 | 56.65 | 50.56 |
Zero-shot cross-lingual: In these experiments, we perform evaluation of Meta4XNLI in the opposite language to that utilised for training. Results are reported in Table 14. XLM-RoBERTa performance exceeds that of mDeBERTa in both languages. Nonetheless, F1 score for EN is almost 20 points lower than ES evaluation results. In addition to the differences in annotation criteria between languages and datasets, another aspect to bear in mind in this scenario could be the number of positive examples present in training sets. As we explained in Section 3.3, Meta4XNLIEN contains a higher number of metaphorical instances, thus models are exposed to a greater variety of examples that can be transferred to ES. While a reduced number of instances in Meta4XNLIES seen during training might hinder the model’s generalization ability, as the low recall and high precision scores show in Appendix Table 28.
@@
In the first setup, we fine-tuned the encoder-only models for the NLI task with the MultiNLI dataset Williams, Nangia, and Bowman (2018). In the case of decoders, we evaluated the models through in-context learning with different prompts. We conducted the evaluation with two different splits for each source dataset in Meta4XNLI: pairs with at least one metaphorical expression and pairs lacking metaphors. From the accuracy scores reported in Tables 15, 16, we can observe certain variability in the results. Encoders’ average performance is close to that of decoders like Qwen2.5-72B-Instruct and gpt-4o. Overall, gpt-4o is the best decoder-only model for EN and ES when evaluated with the CoT prompt.
With respect to encoders, in the majority of cases, XLM-RoBERTa achieves better performance than mDeBERTa. We do observe a tendency for higher results on the sets of pairs without metaphors than on the set with metaphors. The exceptions to this current are the results from XLM-RoBERTa for XNLIdev and XNLItest in ES. In these partitions, the subset of pairs with metaphorical expressions obtained better results, although the difference does not even reach one point.
@@
| average | 83.05 | 82.61 | 83.36 | 83.51 | |
| Llama-3.3-70B-Instruct | Qwen2.5-72B-Instruct | gpt-4o | ||||||||||||
| zero-shot | chain | zero-shot | chain | zero-shot | chain | |||||||||
| EN | ES | EN | ES | EN | ES | EN | ES | EN | ES | EN | ES | |||
| XNLIdev | met | 70.70 | 67.82 | 77.05 | 79.12 | 82.47 | 80.85 | 84.31 | 83.73 | 84.26 | 82.53 | 86.50 | 83.39 | |
| no met | 72.72 | 69.97 | 81.14 | 79.79 | 85.61 | 80.42 | 86.58 | 80.99 | 87.04 | 81.88 | 86.76 | 82.74 | ||
| XNLItest | met | 73.39 | 70.80 | 82.42 | 80.00 | 83.10 | 80.74 | 85.97 | 80.34 | 86.55 | 80.52 | 87.85 | 82.07 | |
| no met | 72.44 | 67.37 | 80.55 | 78.71 | 84.47 | 80.50 | 87.75 | 81.99 | 86.72 | 82.48 | 87.55 | 82.85 | ||
| esXNLI | met | 72.05 | 74.34 | 76.02 | 80.78 | 73.10 | 84.13 | 49.19 | 83.33 | 75.53 | 82.67 | 76.32 | 85.32 | |
| no met | 73.65 | 75.86 | 79.49 | 83.84 | 81.82 | 85.42 | 80.34 | 85.63 | 81.50 | 85.26 | 81.57 | 86.52 | ||
| average | 72.56 | 70.55 | 79.89 | 80.32 | 83.28 | 81.71 | 84.72 | 82.57 | 84.61 | 82.54 | 85.44 | 83.62 | ||
In Table 16, we can also observe the same trend: the subsets without metaphors achieve higher scores. These results are more consistent when evaluated in the language of the original dataset (EN for XNLI and ES for esXNLI), and are more variable in the remaining scenarios. We hypothesize the original language of the dataset might be involved in this performance, since XNLI includes natural utterances of EN that were afterwards manually translated to ES, thus some artifacts might have been introduced during this process Artetxe, Labaka, and Agirre (2020).
The second scenario consisted of fine-tuning the MLMs on two setups: a) with pairs without metaphors and b) pairs with and without metaphors. We performed the evaluation on subsets split by the criterion of metaphor occurrence as well. Results reported in Table 17 show XLM-RoBERTa outperforming mDeBERTa in all contexts. Both models show best results for the NLI task on the examples without metaphors and the lowest performance with pairs that contain metaphorical expressions. This outcome replicates in both languages and experimental setups a) and b).
@@ 6 Error Analysis
In this section, we manually inspected a subset of erroneous cases in order to provide a qualitative insight into the results and with the intention of finding potential explanations of errors and models’ performance for both detection and interpretation tasks.
@@
We selected the predictions from the MLM that obtained highest F1 score from the monolingual experiments in ES for both Meta4XNLI and CoMeta evaluations. We extracted false negatives and false positives and grouped the tokens by their number of occurrences (see Table 18).
| Dataset | Model | Language | FP | FN | Total Errors |
| CoMeta | XLM-RoBERTa | ES | 132 | 175 | 307 |
| Meta4XNLI | mDeBERTa | ES | 137 | 262 | 399 |
| XLM-RoBERTa | EN | 409 | 360 | 769 | |
| Llama-3.1-8B-Instruct | ES | 314 | 313 | 578 | |
| gemma-7b-it | EN | 444 | 551 | 995 |
Within the false positives of CoMeta test set, we find tokens like apoyo (lit. “support”) or tensión (lit. “tension”) that appear more frequently used figuratively than with their literal sense or even only appear in the training set labeled as metaphor. Other wrong predictions are words that appeared in specific domains, such as texts that allude to the pandemic, e.g. ola (lit. “wave”), not detected as metaphorical due to its absence with metaphorical meaning in Meta4XNLI sentences.
Something similar occurs with the misclassified tokens from Meta4XNLI test set. Most errors stem from conventional metaphors, namely gran, abrir, paso, claro (lit. “great”, “to open”, “step”, “clear”) that are regularly used with their metaphorical meaning. The lack of balanced examples might contribute to these predictions. However, we maintained the distribution as is, since our aim is to study the presence and prevalence of metaphor in natural language utterances.
We also conducted an error analysis of the LLMs experiments, focusing on the best-performing model Llama-3.1-8B-Instruct in ES and gemma-7b-it in EN. The observed error patterns show similar trends to those found in the MLMs, particularly in terms of false positive and false negative tokens, in Table 19. Given that Llama-3.1-8B-Instruct achieved a higher F1 score in ES, we also examined cases where the model made correct predictions in ES but failed to do so in EN. Representative examples are provided in Table 20. As illustrated in examples 2 and 5, the EN model makes a prediction, but it does not match the gold standard metaphor. In example 5, for instance, the model correctly identifies “attack” but misses “hard”. In the remaining examples, the model fails to detect the metaphorical token entirely in EN, despite successful identification in the ES counterparts.
| Language | Type | Tokens (Frequency) |
| ES | FN | gran (6), limpio (3), claro (3), gran (2), fusionarán (2), fuerzas (2) |
| FP | gran (4), postura (2), paso (2), llegaron (2), conexiones (2), claro (2) | |
| EN | FN | great (6), forces (4), placed (3), surrounding (2), see (2), scattered (2) |
| FP | great (5), lead (4), high (4), ways (3), support (3), reach (3) |
| # | English Sentence | Spanish Sentence | Gold (EN) | Pred. (EN) |
| 1 | The social relationship of forces is inevitable. | La relación social de fuerzas es algo inevitable. | forces | — |
| 2 | The right-wing parties have provided a link to positively orient the vote of the population. | Los partidos de derechas han facilitado un link para orientar positivamente el voto de la población. | orient | link |
| 3 | If there is a surge of electricity, it’s very dangerous to anyone around. | Si hay una oleada de electricidad, es muy peligroso para cualquiera que esté cerca. | surge | — |
| 4 | It’s hard to install the system because hackers attack it every night. | Es difícil instalar el sistema porque los piratas informáticos lo atacan todas las noches. | attack | hard |
@@
We analysed a subset of 30 errors from each experimental setup, both EN and ES, and evaluation sets. We chose the predictions from XLM-RoBERTa since it is the model that performs better in this task in most scenarios. Although results show that models struggle more to identify the inference relation if there is a metaphorical expression involved in the pair sentences, we do not observe any particular feature within the errors.
A remarkable aspect to highlight is the lower results of esXNLI with respect to XNLIdev and XNLItest in Table 15. It could be motivated by the difference in the text domains since XNLI is an extension of MultiNLI, which maintains the same set of textual domains. While esXNLI is a collection of texts from another set of genres and sources. Regarding EN, some of the errors might derive from the misclassification of some pairs, since annotations were developed on ES text. A metaphorical expression involved in the inference relationship in ES might not be present in its EN version and vice versa. Thus, samples from these two classes, pairs with and without metaphors, should be reexamined in EN and correctly classified for future experimentation.
@@ 7 Conclusions and Future Work
In this work, we present Meta4XNLI, the first cross-lingual parallel dataset in ES and EN labeled for metaphor detection and interpretation framed within the task of NLI. This new resource allows to perform a series of experiments to assess the capabilities of MLMs and LLMs when dealing with this kind of figurative expressions present in natural language utterances.
@@
Regarding the task of metaphor detection, after the annotation process and experimental results, we can conclude the importance of establishing a unified criterion for annotation that is valid for different languages if the aim is to continue researching cross-lingual approaches. In addition, the semi-automatic process of annotation followed for EN shows that automatic labeling of cross-lingual metaphor is far from trivial. Metaphorical expressions are language and culture-dependent. Moreover, the translation of the data introduces a new layer in which metaphorical expressions can either be lost from the translation of the source to the target language, or can be introduced in the target language by means of the translator, either human or automatic. Further work to explore automatic annotation methodologies would be of considerable value in reducing the demanding workload and effort of manual labeling in more than one language.
@@
With respect to results, the purpose of our experiments is not to outperform the state-of-the-art results, but to analyze LMs’ capabilities in processing metaphorical language in a multilingual and crosslingual setting. Best results are obtained when Meta4XNLI in both languages is used for training. The augmentation of the training set size and the parallel annotations might boost this performance. Cross-domain and monolingual experiments show how the lack of consistency in the annotation criteria affects the performance of models. This can be observed as well in the zero-shot cross-lingual setup, although the scenario of training in EN and evaluating in ES shows competitive performance. It should be noted that the EN set contains a larger number of instances annotated as metaphorical in the training set. In addition, the in-vocabulary and out-of-vocabulary evaluation points to some kind of bias in the learning process. This could stem from the fact that the majority of the metaphor instances are conventional or due to lexical memorization Levy et al. (2015); Boisson, Espinosa-Anke, and Camacho-Collados (2023). Future research on this line of work should be carried out to clarify this issue.
@@
Regarding metaphor interpretation, we evaluated the ability of MLMs and LLMs to understand metaphorical expressions framed within the NLI task. We provide parallel annotations at the premise-hypothesis pair level that mark whether the presence of metaphorical expressions is relevant for the inference relationship. We exploited this information to conduct our experiments. From the reported results, we can observe a tendency of the models to perform lower with pairs that contain at least one metaphorical expression. However, this trend breaks when evaluating the datasets in their translated version. We presume the translation process might induce biases in metaphor occurrence and the “naturalness” of the sentences. Similarly to metaphor detection, future work to analyse the impact of translation in the development of metaphor parallel resources should be explored for the task of metaphor interpretation, as well as additional experimentation from a multilingual perspective.
In summary, our work provides a high-quality, cross-lingual and parallel resource with aligned annotations for detection and interpretation over the same text. Our new dataset not only facilitates systematic evaluation of model performance across languages but also serves as a starting point for future research in metaphor detection, interpretation, and metaphor transfer across languages. By tackling both annotation and empirical challenges, we lay the groundwork for more accurate and critical assessments of how language models handle metaphor and meaning across languages.
@@ 8 Limitations
Metaphor annotation is an inherently subjective task. This variance in annotations is reflected in Meta4XNLIEN, due to the different criteria employed through the annotation process. Labels in this language should be updated and further revised to improve their quality. Disagreement and subjectivity could be counterbalanced by more annotation iterations and a larger number of annotators, in order to develop more consistent and reliable labeled data. However, this constitutes an arduous and costly process. Data augmentation and semi-automatic methods could be exploited to create larger datasets with similar characteristics to the one we present and extend it to more languages, since most corpora available for metaphor processing are of reduced size and limited to a narrow set of languages. The existence of parallel resources in multiple languages other than EN that reflect cultural and real-world knowledge nuances is of great importance to continue researching such a complex phenomenon as figurative language, specifically metaphors.
Data Splits Statistics
@@
Inter-Annotator Agreement Labeling Process
Detection Experiments Results
@@
| Llama-3.1-8B-Instruct | Qwen2.5-7B-Instruct | gemma-7b-it | ||||||||
| Train | Test | F1 | Prec. | Rec. | F1 | Prec. | Rec. | F1 | Prec. | Rec. |
| M4XES | M4XES | 59.15 | 59.11 | 59.19 | 56.76 | 65.98 | 49.80 | 58.06 | 65.47 | 52.15 |
| CoMeta | 50.06 | 53.56 | 46.99 | 44.26 | 53.07 | 37.96 | 45.32 | 52.60 | 39.81 | |
| M4XEN | M4XEN | 50.41 | 53.68 | 47.52 | 47.66 | 44.79 | 50.92 | 52.61 | 55.50 | 50.00 |
| VUAM | 26.78 | 61.68 | 17.10 | 26.44 | 59.72 | 16.98 | 26.72 | 70.19 | 16.50 | |
| Meta4XNLIES | Meta4XNLIEN | |||||
| F1 | Prec. | Rec. | F1 | Prec. | Rec. | |
| mDeBERTa | 68.70 1.27 | 72.78 3.45 | 65.29 3.77 | 60.80 1.05 | 65.03 2.08 | 57.23 3.10 |
| XLM-RoBERTa | 66.01 0.55 | 70.35 4.42 | 62.53 3.93 | 60.14 1.32 | 64.97 2.81 | 56.15 3.04 |
| Llama-3.1-8B-Instruct | 61.46 | 61.90 | 61.02 | 55.26 | 53.33 | 57.34 |
| Qwen2.5-7B-Instruct | 58.72 | 59.31 | 58.15 | 52.62 | 47.56 | 58.90 |
| gemma-7b-it | 57.02 | 64.89 | 50.85 | 54.60 | 57.79 | 51.74 |
@@
Prompts for Inference on Interpretation
| Zero-shot |
| Say which is the inference relationship between these two sentences. Please, answer only with one word between ’entailment’, ’neutral ’or ’contradiction’. |
| {Premise} -> {Hypothesis}: |
| Chain-of-thought |
| You are an expert linguist and your task is to annotate sentences for the task of Natural Language Inference. This task consists in determining if a first sentence (premise) entails, contradicts or does not entail nor contradict the second sentence (hypothesis). Please, answer with one of the following labels: ’entailment’, ’contradiction’ or ’neutral’. |
| Here you have a few examples: |
| Premise: I am an open person. |
| Hypothesis: I am friendly. |
| Answer: entailment |
| Premise: My heart is broken. |
| Hypothesis: I am happy. |
| Answer: contradiction. |
| Premise: You are a close friend. |
| Hypothesis: I like your eyes. |
| Answer: neutral. |
| {Premise}: |
| {Hypothesis}: |
| {Answer}: |
Acknowledgements.
This work has been supported by the HiTZ center and the Basque Government (Research group funding IT-1805-22). Elisa Sanchez-Bayona is funded by the UPV/EHU PIF20/139 grant. We also thank the funding from the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and ERDF A way of making Europe; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR.References
[…]
- Dubey et al. (2024) Dubey, Abhimanyu, Abhinav Jauhri, Abhinav Pandey, et al. 2024. The llama 3 herd of models.
- García-Ferrero et al. (2024) García-Ferrero, Iker, Rodrigo Agerri, Aitziber Atutxa Salazar, Elena Cabrio, Iker de la Iglesia, Alberto Lavelli, Bernardo Magnini, Benjamin Molinet, Johana Ramirez-Romero, German Rigau, Jose Maria Villa-Gonzalez, Serena Villata, and Andrea Zaninello. 2024. MedMT5: An open-source multilingual text-to-text LLM for the medical domain. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 11165–11177, ELRA and ICCL, Torino, Italia.
- Jang et al. (2014) Jang, Hyeju, Mario Piergallini, Miaomiao Wen, and Carolyn Rosé. 2014. Conversational metaphors in use: Exploring the contrast between technical and everyday notions of metaphor. In Proceedings of the Second Workshop on Metaphor in NLP, pages 1–10, Association for Computational Linguistics, Baltimore, MD.
- Kesarwani et al. (2017) Kesarwani, Vaibhav, Diana Inkpen, Stan Szpakowicz, and Chris Tanasescu. 2017. Metaphor detection in a poetry corpus. In Proceedings of the Joint SIGHUM Workshop on Computational Linguistics for Cultural Heritage, Social Sciences, Humanities and Literature, pages 1–9, Association for Computational Linguistics, Vancouver, Canada.
- Kwon et al. (2023) Kwon, Woosuk, Zhuohan Li, Siyuan Zhuang, Ying Sheng, Lianmin Zheng, Cody Hao Yu, Joseph E. Gonzalez, Hao Zhang, and Ion Stoica. 2023. Efficient memory management for large language model serving with pagedattention. In Proceedings of the ACM SIGOPS 29th Symposium on Operating Systems Principles.
- Leong, Chee Wee and Beigman Klebanov, Beata and Shutova, Ekaterina(2018) (Ben) Leong, Chee Wee (Ben) and Beigman Klebanov, Beata and Shutova, Ekaterina. 2018. A Report on the 2018 VUA Metaphor Detection Shared Task. In Proceedings of the Workshop on Figurative Language Processing, pages 56–66, Association for Computational Linguistics, New Orleans, Louisiana.
- Naik et al. (2018) Naik, Aakanksha, Abhilasha Ravichander, Norman Sadeh, Carolyn Rose, and Graham Neubig. 2018. Stress test evaluation for natural language inference. In Proceedings of the 27th International Conference on Computational Linguistics, pages 2340–2353, Association for Computational Linguistics, Santa Fe, New Mexico, USA.
- Sánchez-Montero et al. (2025) Sánchez-Montero, Alec, Gemma Bel-Enguix, Sergio-Luis Ojeda-Trueba, and Gerardo Sierra. 2025. Disagreement in metaphor annotation of Mexican Spanish science tweets. In Proceedings of Context and Meaning: Navigating Disagreements in NLP Annotation, pages 155–164, International Committee on Computational Linguistics, Abu Dhabi, UAE.
- Team et al. (2024) Team, Gemma, Thomas Mesnard, Cassidy Hardin, Robert Dadashi, Surya Bhupatiraju, et al. 2024. Gemma: Open models based on gemini research and technology.
- Team (2024) Team, Qwen. 2024. Qwen2.5: A party of foundation models.
- Tkachenko et al. (2020-2025) Tkachenko, Maxim, Mikhail Malyuk, Andrey Holmanyuk, and Nikolai Liubimov. 2020-2025. Label Studio: Data labeling software. Open source software available from https://github.com/HumanSignal/label-studio.
- Wang et al. (2024) Wang, Shun, Ge Zhang, Han Wu, Tyler Loakman, Wenhao Huang, and Chenghua Lin. 2024. MMTE: Corpus and metrics for evaluating machine translation quality of metaphorical language. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages 11343–11358, Association for Computational Linguistics, Miami, Florida, USA.